OpenStudy (anonymous):

2 square feet is same as how many square inches

OpenStudy (anonymous):

give answer

OpenStudy (anonymous):

144 in one foot so multiple 144 by 2, and that is....288

OpenStudy (anonymous):

u r girl or boy

OpenStudy (anonymous):

gurl....????

myininaya (myininaya):

12 inches=1 foot 12 square inches = 1 square foot 2*12 square inches=2*1 square foot 24 square inches = 2square feet

OpenStudy (anonymous):

good night friends byeeeeeeeeeeeeeeeeeeeeeeeeeeee

OpenStudy (anonymous):

uuummmmmm bii

OpenStudy (anonymous):

what that mean

OpenStudy (anonymous):

randichandi

myininaya (myininaya):

i think it means bye

OpenStudy (anonymous):

ys

OpenStudy (anonymous):

ok luv u

OpenStudy (anonymous):

luv2u2

OpenStudy (anonymous):

u boi or gurl?

OpenStudy (anonymous):

u r girl

OpenStudy (anonymous):

u?

OpenStudy (anonymous):

i am boy

OpenStudy (anonymous):

o kkool